### v1

# From attention to profit: quantitative trading strategy based on transformer

Abstract
In traditional quantitative trading practice, navigating the complicated and dynamic financial market presents a persistent challenge. Former machine learning approaches have struggled to fully capture various market variables, often ignore long-term information and fail to catch up with essential signals that may lead the profit. This paper introduces an enhanced transformer architecture and designs a novel factor based on the model. By transfer learning from sentiment analysis, the proposed model not only exploits its original inherent advantages in capturing long-range dependencies and modelling complex data relationships but is also able to solve tasks with numerical inputs and accurately forecast future returns over a period. This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2019. The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies with lower turnover rates and a more robust half-life period. Notably, the model’s innovative use transformer to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies.
keywords:
Quantitative finance, Transformer, Stock selection, Market sentiment
1 Introduction
The goal of stock trading is to optimize the return on the investment in the capital market according to the process of buying or selling one or more companies’ shares. Traders obtain profit when the difference is generated by the fluctuation of stock spice. However, stocks are influenced by a large number of factors, which constitute a complex system and increase the difficulty for people to make a profit. The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling [1]. So, many strategies and tools were built, which corresponded to the development of the capital market and quantitative strategies have played important roles among them.
Figure 1: The Quantitative Trading Architecture with Factors
Some traditional quantitative tools, such as Markowitz Portfolio [2] and Capital Asset Pricing Model(CAPM) [3], focuses mainly on the fundamental analysis. In other words, these strategies aim to make a profit by simple calculation and analysis. However, depending on the development of computer science, more quantitative methods and tools were introduced. Within these methods, strategy with factors has attracted people’s attention. In 1993, Fama and French [4] introduced their Fama French Three Factor Model (FF3), which is one of the most influential models in quantitative trading. In 2015, Fama and French renewed their model, the Five-Factor Asset Pricing Model (FF5) [5] was introduced. Besides this classical theory, other trading strategies have been published for decades.
Quantitative trading with factors typically manifests in two primary strategies, which are shown in Fg.1. The first approach involves the computation of stock factor values. Based on these calculated values, stocks are ranked to establish a pool. Once this pool is established, assets are held for a predetermined period. Adjustments to the portfolio are then made at specific time intervals, ensuring alignment with evolving market conditions and factor readings. The second method employs a fixed pool of stocks, wherein factors guide the derivation of long/short signals. Traders can execute corresponding actions when they receive the signals from factors, allowing for a dynamic response to market fluctuations based on factor insights.
In recent years, Machine Learning (ML) stands out as an instrumental tool in enhancing trading algorithms and decision-making processes. Its core principle, enabling systems to learn from and make decisions based on data, lends itself particularly well to the vast and dynamic landscapes of stock markets. Some traditional literature exploits machine learning algorithms based on manual indicators [6], but the hypothesis about the random walk may be less useful in the task of catching non-stationary series. With the rise of different architectures, quantitative researchers found some novel approaches to identifying intricate patterns and factors in trading data, which opened a new field in quantitative trading [7].
Although there are several previous experiments about ML factors in the quantitative finance field, the research still faces two concerns. Firstly, in the field of NLP, it is necessary to convert words in text to word vectors through word embeddings to serve as inputs. However, in the financial dataset, there are categorical data such as industry types, as well as numerical data like price fluctuation, turnover rate, and financial indicators. If the input comprises only categorical data, the time series can be treated as a sentence [8]. In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
For the second concern, in NLP tasks, most of them can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition. To utilize existing outputs, decoders in NLP tasks sequentially output samples and use masking operations to handle input sequences during training. However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
To address these concerns, we propose an approach that combines time series stock data with a numerical-adapted transformer to establish a factor and put it into back-tests. This approach allows us to use transformer model embedding numerical input directly, which refers to a method similar to sentiment analysis. The structure of the paper is designed as follows. Section 2 discussed the previous quantitative financial works mainly based on machine learning. Section 3 and 4 introduce the establishment of the model. A factor based on the model will be trained and back-tested. For the practical test, we collected data from more than 4,600 stocks in the previous 14 years (from 2010 to 2023) from the data-collect platform. To comprehensively test the ability of the factor, we divided of data by different frequencies (monthly, weekly, and daily). Finally, the result of the back-tests including the comparison between the transformer-factor and others and the insights gained from such comparative analysis will be discussed in the section 4 and 5.
Figure 2: Example for stock prediction
2 Preliminaries
This section briefly introduces the related works about the paper and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
2.1 Stock data with market sentiment
Stock data can be considered that can influence the future trend of this stock in some cases [9]. Previous works have discussed the correlation between investor sentiment and stock market metrics like turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns [10, 11]. This relationship isn’t uniform but varies with changes in investor sentiment. Research also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions. This link between sentiment and market performance is further affirmed by studies focusing on factors like the turnover ratio, which are seen as reflective of market liquidity and investor behavior [12]. These findings underscore the mutual relationship between investor sentiment and market performance, demonstrating how psychological factors can drive market dynamics.
2.2 Applications for ML methods in trading
Machine learning exploits a range of algorithms and statistical techniques used for tasks such as regression, clustering, and classification. In past decades, there have been methods that have been significantly increasingly applied in stock due to their ability to process vast amounts of data and make predictions, offering a potential advantage in the financial markets. At the same time, the finance sector, particularly quantitative trading, has started to be aware of the potential of deep learning models for predicting stock movements, portfolio optimization, and risks.
2.2.1 Support Vector Machines (SVM)
SVM was widely used in quantitative trading, which can manage high-dimensional spaces and intricacies found in datasets. In 2003, SVM tried to forecast financial time series [13], especially with the lens focused on the South Korean stock market. This work exhibited the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting. Similar to the previous one, [14] showcased SVM’s prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments. [15], found SVM’s adaptability and precision in financial tasks like portfolio and quantitative trading.
However, SVM also shows some disadvantages in quantitative trading. In 2009, SVM was underscored for the accuracy of overfitting [16]. When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, causing the strategy ineffective on unseen data. In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptive [17].
2.2.2 Long Short-Term Memory (LSTM)
LSTM was designed to capture long-term dependencies in sequential data according to their unique architecture [18]. [19] deployed LSTM with stacked autoencoders to forecast financial time series. LSTM shows the ability to accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models. In 2018, [20] tested the efficacy of LSTM for price predictions. They observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over some conventional models. LSTM’s capability to extract features from time series was explored by [21]. They combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy. For predicting stock prices, [22] harnessed the power of LSTM to unearth multi-frequency trading patterns. Their research underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
2.2.3 GRU
Gated Recurrent Units (GRU) [23] have emerged as a useful method in the quest for accurate financial forecasting. GRU can capture long-term dependencies in time series data and has seen increasing popularity in the world of quantitative trading. [19] introduced a comprehensive deep learning framework for financial time series, which Incorporated GRUs alongside LSTM and stacked autoencoders. Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs. [24] centralized on LSTM networks, provided comparative insights into GRUs. However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints. What’s more, [25] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [26]. The strategy combined CNNs and GRUs, also received advantages among them, creating a formidable solution for stock price prediction.
2.3 Transformer in quantitative trading
2.3.1 Introduction for transformer
Transformer [27], which was introduced in 2017, has received high attention in the NLP field. According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models. The transformer-based models like GPT [28] and BERT [29] are the basis of the specific field of application. In the application field, transformer-based models like ChatGPT [30], FinBERT [31, 32] and BloomberGPT [33] show advantages and ability in solving different types of problems in different filed.
Similar to other neutral sequence transduction models, transformer also has an encoder-decoder structure. The model takes input as a sequence of context which could be natural language sentences and generates output with texts depending on the input, like a translation according to the input or a predicted sentence. Transformer using stacks which are called self-attention and point-wise, which connect layers in the architecture for encoder and decoder. For these encoder and decoder layers, the self-attention method plays an important role in matrix calculation. The linear and softmax functions are used in the dot-product attention module. In the vanilla transformer model, there are multi-head attention layers are used and the result from each layer will be outputted after going through an additional &\&& normalize layer. The above-mentioned steps covered most of the progress in the vanilla transformer.
2.3.2 Previous transformer-based strategy
Transformer has achieved superior performance on long-term dependencies modeling. In the quantitative trading domain, some previous researchers tried to test the potential of transformer. [34] firstly tried to exploit Transformer on trading sequences to classify stock price movements.
Besides using a traditional transformer, other innovative methods of transformer are worth mentioning. [35] devoted to optimizing the efficiency of time complexity and memory usage of Transformer on extremely long time series by informer. [36] leveraged the advantages of CNNs and Transformers to model short-term and long-term dependencies in financial time series. They showed the advantage of the proposed approach on intraday stock price prediction of S&P 500 constituents.
Similar to previous work, [1] innovated the structure of a Transformer, which is Adaptive Long-Short Pattern Transformer (ALSP-TF). Their model is structurally innovated for hierarchical representation and interaction of stock price series at different context scales. With the help of a learnable function, they make the self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching. Finally, they obtained more than 10% of annual return on average.
However, there are some disadvantages of transformer. [37] mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights. As fluctuations of stocks are conditioned on composite signals over manifold periods, lacking pattern-wise interaction hinders the adequate discrimination of stock tendency and is susceptible to noise points. On the other hand, [38] claimed that the basic query-key matching paradigm is position agnostic. Although position embedding is inserted into the sequential inputs, it may not be optimal because of the inability to reveal precise distances.
3 Methodology
This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, model enhancement and output processing.
3.1 Framework overview
Fig. 3 illustrates the overview of the structure. There are three major parts:
(i) Data Initialization to align the time series of stocks into a regular matrix;
(ii) Embedding the training dataset and training them by transformer which replaces the word-embedding layer to linear layer and deletes the Mask(opt.) layer;
(iii) Lastly, using the trained model to predict the possible trend of stocks.
Figure 3: Overview of the work
3.2 Problem formulation
Inputs
Given the candidate stock set $S^{t}$ for N stocks on the trading timestamp $t$:
$S^{t}=\{s_{1}^{t},s_{2}^{t}\ldots s_{N}^{t}\}$
For a stock $s_{k}^{t}$ in stock set $S^{t}$, where $k\in N$, it details a sequence for a historical 1-time step (such as one month), and the dimension is 2:
${\chi}^{t}_{k}=[x_{k}^{t-20},x_{k}^{t-19}\ldots x_{k}^{t}]\in\mathbb{R}^{20\times 2}$
In each $x_{k}^{t}$, there are two features. The first one is the accumulated daily profit rate during the time step, where $p_{i}$ and $p_{i-1}$ is the close price of the $i^{th}$ and ${i-1}^{th}$ day, respectively.
$r^{t}=\sum r_{i}=\sum\frac{p_{i}-p_{i-1}}{p_{i-1}},\quad\sigma^{t}=\sum\sigma_{i}$ (1)
Another is the accumulated daily turnover rate $\sigma^{t}$ during the time step and the calculation method of $\sigma^{t}$ is shown above. So each $\chi^{t}_{k}$ can be represented as:
$\chi^{t}_{k}=\left[\begin{matrix}r^{t-20}&r^{t-19}&\ldots&r^{t}\\ \sigma^{t-20}&\sigma^{t-19}&\ldots&\sigma^{t}\\ \end{matrix}\right]\in\mathbb{R}^{20\times 2}$
Outputs
For a stock set $S^{t}$ on the trading timestamp $t$, each stock $S^{t}_{i}$ has a profit rate $r_{i}^{t+1}$ on the trading timestamp $t+1$, where the calculation method is same as the method of $r^{t}$ in the sequence of inputs shown in Eq.(1). Then the value in the list of next-time stamp’s profit for $N$ stocks in the timestamps $t+1$ can be represented as:
$P^{t+1}=\{r_{1}^{t+1},r_{2}^{t+1}\ldots r_{N}^{t+1}\}$
And the list will be arranged as $n$ parts and each part has an equal proportion $\frac{1}{n}$ ($n$ should be an odd number, such as 3,5,7). Take $n=5$ as an example. For those ranked values in the first, middle and last $\frac{1}{5}=20\%$, the output of the stock $s_{k}$ records as below, respectively:
$y^{t}_{k}=[1,0,0]^{T},\quad y^{t}_{k}=[0,1,0]^{T},\quad y^{t}_{k}=[0,0,1]^{T}$
3.3 Embedding initialization
Normalization
The next step is the normalize the input data for each time step by zero-mean normalization (Z-Score), the normalization equation is shown in Eq.(2).
$\tilde{x}^{t}_{k,i}$ represents the normalized value at $i^{th}$ day of stock $k$ in time $t$.
$E[x^{t}_{i}]$ and $std[x^{t}_{i}]$ is the mean and standard of all the value of $x^{t}_{i}$, respectively.
$\tilde{x}^{t}_{k,i}=\frac{x^{t}_{k,i}-E[x^{t}_{i}]}{std[x^{t}_{i}]}$ (2)
In this way, the input consequences towards zero mean and unit variance, which aims to reduce the influence from different time points and allow different features to be comparable with each other [39].
3.4 Transformer encoder
Then a transformer encoder structure will be built in this subsection. The representation hierarchy consists of $L$ blocks of multi-head self-attention layers. Taken initialized stock embedding sequences $\textbf{P}\in\mathbb{R}^{N\times 20\times 2}$ as input, and the canonical self-attention in [27] performs information exchange between every time points for each stock $s_{i}$.
To process both categorical and numerical data, the word embedding layer is replaced by a standard linear layer, utilizing linear transformations to substitute the process of word embedding. On the other side, in stock prediction, we aim to accurately forecast the return for a future period, thus the model’s output is generally a single value or the probability of price increase or decrease. Therefore, the decoder is simplified by removing the autoregressive prediction mechanism (in the decoder) and the masking operations.
For each $\textbf{P}_{i}\in\mathbb{R}^{20\times 2}$, transforms into query, key, and value metrics, respectively:
$\textbf{Q}_{i,h}=\textbf{E}_{i}\textbf{W}^{Q}_{h},\quad\textbf{K}_{i,h}=\textbf{E}_{i}\textbf{W}^{K}_{h},\quad\textbf{V}_{i,h}=\textbf{E}_{i}\textbf{W}^{V}_{h}$ (3)
where $h=1,\ldots,H$ is the head index, as well as $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$, $\textbf{W}^{Q}_{h}$ $\in\mathbb{R}^{2\times d_{f}}$. Lastly, the final layer is represented by the concatenation of all attention heads.
$\textbf{F}_{i}=Multihead(\textbf{Q}_{i,h},\textbf{K}_{i,h},\textbf{V}_{i,h})=||^{h=H}_{h=1}Softmax(\textbf{Q}_{i,h},\textbf{K}_{i,h}^{T}/\sqrt{d_{f}})\textbf{V}_{i,h})$ (4)
Just like Eq.(4) shows, ||||| | represents the concatenation operator, and $d_{f}$ is the dimension of projected feature space. In the end, all input consequences are represented in the form of:
$\textbf{F}=[\textbf{F}_{1};\textbf{F}_{2};\ldots\textbf{F}_{N}]\in\mathbb{R}^{N\times 20\times Hd_{f}}$
which is followed by feed-forward layers.
According to Sec.3.2, the embedding output is denoted as:
$\textbf{Y}^{t}=\{y^{t}_{1},y^{t}_{2},\ldots y^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$
3.5 Output process
Eventually, the output is represented:
$\textbf{Z}^{t}=\{z^{t}_{1},z^{t}_{2},\ldots z^{t}_{N}\}\in\mathbb{R}^{N\times 1\times n}$ (5)
The output will be activated by the $softmax$ function. And for each $z^{t}_{i}$, there are:
$z^{t}_{i}=[z^{t}_{i,1},\ldots z^{t}_{i,m},\ldots z^{t}_{i,n}]\in\mathbb{R}^{1\times n}$
the output of probability can be represented as (m is the middle number from 1 to n):
$Y^{t}_{o}=[y^{t}_{o,1},\ldots y^{t}_{o,m},\ldots y^{t}_{o,n}]^{T}=[\frac{e^{z_{1}}}{\sum_{i=1}^{n}e^{z_{i}}},\dots\frac{e^{z_{m}}}{\sum_{i=1}^{n}e^{z_{i}}},\ldots\frac{e^{z_{n}}}{\sum_{i=1}e^{z_{i}}}]$ (6)
In Eq.(6), $y^{t}_{o,i}$ are larger than 0 and smaller than 1 as well as the sum of all the items in $Y^{t}_{o}$ is 1. So the output means the probability of the stock’s raising, flatting, or downing in the next timestamp.
3.6 Prediction
The Mean Squared Error Loss (MSELoss) will be used to quantify the loss. MSE Loss is a widely used metric for assessing the discrepancy between a model’s predictions and the actual values. It calculates the average of the squares of the differences between the predicted value and the actual prices. This is formally represented as:
$MSELoss=\frac{1}{n}\sum^{n}_{i=1}(y_{i}-\hat{y}_{i})^{2}$ (7)
In Eq.(7), $y_{i}$ is the actual value of the $i^{th}$ equation, $\hat{y}_{i}$ is the model’s prediction for the same, and n represents the total number of samples in the inputs. During each epoch, the MSE Loss is computed to guide the optimization process, to minimize this value over successive iterations. A lower MSE indicates a closer alignment between the model’s predictions and the actual stock prices, signifying an improvement in the model’s learning and its ability to generalize from the training data. This loss function is particularly useful in regression tasks, as it not only penalizes the model for inaccurate predictions but also emphasizes reducing larger errors, thereby ensuring a robust learning process.
4 Experiments
Based on the methodology in Section 3, in this section, the set of experiments is introduced, including the data resource, implementation details, trading strategy and metrics.
4.1 Dataset
The training data of the transformer factor comes from the Chinese exchange market. The dataset is collected from the Shanghai Stock Exchange (SHSE) and Shenzhen Stock Exchange (SZSE), which contains 4601 stocks that are listed or had been listed from January 2010 to May 2023. And all these data come from AKShare [^1] https://akshare.akfamily.xyz/ and Tushare [^2] https://www.tushare.pro/, which is one of the quantitative finance terminals. The train period is from January 2010 to December 2019 and the test period starts from January 2020.
Adjustment closing price is used for the stock training price [40]. Adjustment closing price is essential in stock market analysis, particularly when analyzing historical data for long-term trends and patterns. This price takes into account factors like dividends, stock splits, and other corporate actions that can affect a stock’s price over time [41]. By adjusting for these events, the adjusted close price provides a more accurate picture of a stock’s value and performance, which is more persuadable in quantitative financial back-test.
4.2 Timestamp
To test the performance of the factor better, three kinds of frequency are chosen, which are stock data monthly, weekly, and daily, respectively.
In the first experiments, a timestamp was set as one month. So for a stock $s_{i}$ from stock set $S^{t}$, it details a feature sequence in which each item in the sequence contains the accumulated profit and accumulated turnover rate in a month. Both the accumulated profit and accumulated turnover rate are the sum of the trading day’s data in that month. In some cases, if one stock was first listed on the market or resumed trading in one month, the available data may not cover whole trading days, the data of the stock from that from would still be recorded. However, if there is a whole month that the stock did not have the trading record (perhaps stopped trading or not listed in that month), this month’s data will be recorded as ”NaN”. In the sequence $\chi^{t}_{k}$, the sequence will not be used to train the model when there are missing values in the sequence.
Within the first experiments, three sub-experiments were set. For the sub-experiments, the input of them are same. For output data, outputs with different dimensions and different lengths but the same dimension are used. $n=5$ and $n=7$ are used in this experiments. In the group of $n=5$, the first one deleted the data with output like $[0,0,0]^{T}$, and the second one used the origin dataset (zero-output).
Most of the accumulated profit values in different years are in the range $(-10\%,10\%)$. There are some values point outside the range $(-100\%,100\%)$ that are shown in it because these values are too large (small), such as the stock ”002961.SZ” first listed at SZSE on September $4^{th}$, 2019 and its accumulated profit rate on that month is 184.66%. These stocks received unusual performance mainly because they were first listed on the market, resumed trading, or reported great positive/negative news.
Most of the accumulated turnover rate values are in the range $(0,100\%)$. There are some values outside the range $(0,200\%)$ that are shown in it because these values are too large, such as the accumulated turnover rate in July 2015 of ”002596.SZ” is $777.67\%$, this was because the actual controller abundantly increased his holdings of the company and the company announced that they received findings by shareholders. Similar to the example, the situation may be caused by great positive or negative information.
Outliers of both accumulated profit and accumulated turnover rate will not be removed from the dataset these situations may happen in the future and they are expected to be predicted, though these incidents rarely happen.
In the second and third experiments, the inputs are in weekly and daily frequency, respectively. The sub-experiments under different groups are similar to the first one. All the accumulated number of parameters, the number of trained sections (such as 100 months from 2010 to 2019), and dimension are shown in the Table 1.
Table 1: Detailed information of experiments
Strategy Frequency Output Dim Training samples Section zero-output
Month_1 Monthly 3 85,490 100 w/o
Month_2 Monthly 3 142,409 100 w/
Month_3 Monthly 5 142,409 100 w/
Week_1 Weekly 3 455,157 466 w/o
Week_2 Weekly 3 758,300 466 w/
Week_3 Weekly 5 758,300 466 w/
Day_1 Daily 3 3,586,435 2,420 w/o
Day_2 Daily 3 5,140,279 2,420 w/
Day_3 Daily 5 5,140,279 2,420 w/
4.3 Implementation Details
The training process and model are implemented with PyTorch [42]. The parameter of the model is searched by gird search, which is reliable in low dimensional spaces and simple to implement [43]. As a result, in the situation of input dimension is 2, the dimension of hidden feature space $d_{f}$ is 16. The number of multi-head attention modules is 16, and the number of layers of encoder and decoder is 6. The model was trained on an NVIDIA GeForce RTX 2080 GPU and NVIDIA A100 Tensor Core GPU by Adam optimizer [44] for 50 epochs, the learning rate is 0.001, and the batch size is 64.
4.4 Trading strategy
A trading strategy is used for the model and Algorithm (1) shows the algorithm of strategy. Before the first trade date of the timestamp $t$, all sequence $\chi^{t}_{i,k}$ from stock set $S^{t}$ will be put in the model and obtain the list of the output. Then the stocks will be ranked according to the first element of the output and the first $\frac{1}{n}\%$ stocks will be added to the stock pool. If the stock has already been in the stock pool on the last timestamp, it will keep holding; if the stock is in the predicted pool but not in the previous pool, it will be bought in with the same proportion of the whole account. Stocks that are not in the predicted pool will be sold out. The same method will run repeated during the rest period.
The backtest starts from January 2020, in other words, the result of the sequences from May 2018 to December 2019 will be used as the first stock pool to trade.
Algorithm 1 Trading Strategy
1:The sequence $x^{t}_{i,k}$ for each stock $s_{i}$ in the set $S^{t}$ and the amount of cash.
2:run_monthly(trade, monthday=1, time=‘open’)
3:function obtain_stockpool($x^{t}_{i,k}$):
4:     stockpool $=$ sort(model($x^{t}_{k}$), key=lambda x : x[1], reverse=True)
5:     return stockpool
6:end function
7:function trade
8:     stockpool $=$ obtain_stockpool($x^{t}_{i,k}$)
9:     for stock in previous_stockpool do
10:         if stock not in stockpool then
11:              order(stock, 0) $\triangleright$ sellout
12:         end if
13:     end for
14:     for stock in stockpool do
15:         if stock not in previous_stockpool then
16:              order(stock, amount / len(stockpool)) $\triangleright$ buy
17:         end if
18:     end for
19:end function
4.5 Metrics
Formally, the Sharpe Ratio (SR) and the $\alpha$ ratio will be used to test the performance of the strategy. The SR [3] is a measure of risk-adjusted return which can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8), where $R_{p}$ is the return of the portfolio, $R_{f}$ is the risk-free rate:
$SR=\frac{E[R_{p}]-R_{f}}{std[R_{p}]}$ (8)
Alpha represents the excess return of an investment portfolio relative to its benchmark. It reflects the stock selection skills of an investment. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment. Specifically, alpha is the excess return of the actual portfolio return over its expected theoretical return. Eq.(9) shows the calculation method of alpha rate, where $E(R_{m})$ is the market return and $\beta$ is the correlation between the portfolio and systemic risk.
$Alpha=(R_{p}-r_{f})-\beta(E[R_{m}]-r_{f})$ (9)
Besides the SR ratio and alpha rate, the annual return (AR) of the stock, the annual excess return (AER), the average turnover rate (TR) of the portfolio, and the win rate (WR) will also shown. The excess returns are returns achieved above and beyond the return of a proxy and the CSI 300 index is used as the basis when calculating the excess returns. The turnover rate is the average percentage of the portfolio adjusted daily in terms of market. The win rate is the percentage of trade days where the portfolio encounters a positive return overall trade days.
5 Result and discussion
Building upon the analysis of individual strategies, this section presents an evaluation and discussion of the experimental outcomes. The discussion focuses on the comparative effectiveness of monthly, weekly, and daily strategies. The results underscore the nuanced relationship between trading frequency and market performance, offering valuable insights for the development of robust quantitative trading strategies.
5.1 Overall performance
With the starting amount of RMB 5,000,000, the results of different strategies are obtained. The result of the experiment is shown in Table 2. For the monthly group, Month_1 shows an annual return of 17.35% and an annual excess return of 19.43%, indicating a robust performance over that period. Conversely, Month_3 shows a diminished annual return and annual excess return of 7.37% and 9.91% respectively. The turnover rate is notably high in the second month at 51.69%. The win rate ranges from 43.1% to 57.3%, with the first month showing the highest win rate. The Sharpe Ratio and Alpha, vary across the strategies, with some periods showing negative values, suggesting underperformance relative to the risk taken, while others show positive values.
Table 2: Result of experiments
Strategy AR AER TR WR SR Alpha
Month_1 17.35% 19.43% 26.09% 57.8% 0.915 0.162
Month_2 9.91% 13.86% 51.69% 49.3% 0.289 0.102
Month_3 7.37% 9.91% 32.33% 51.6% 0.246 0.064
Week_1 -0.83% 1.31% 7.13% 46.4% -0.236 -0.03
Week_2 7.49% 10.81% 1.18% 49.2% 0.160 0.085
Week_3 12.3% 12.73% 1.39% 54.4% 0.372 0.116
Day_1 7.89% 11.4% 6.71% 43.1% 0.181 0.090
Day_2 10.23% 10.94% 6.51% 44.4% 0.279 0.097
Day_3 9.81% 10.03% 5.57% 44.6% 0.281 0.092
CSI300 1.77% \ \ \ \ \
In the weekly strategy results, Week_3 showed a remarkable annual return of 12.3% and an annual excess return of 12.73%, while Week_1 displayed a negative annual return. For the daily strategies, Day_2 had a return of 10.23% and an annual excess return of 10.94%. The turnover rate across the daily strategies showed lesser variation than the monthly or weekly, indicating a more consistent trading frequency. The win rate for the daily strategies remained fairly stable, hovering around the mid-40% mark. The Sharpe Ratio and Alpha for weekly and daily strategies show a mix of positive and negative values, reflecting the fluctuating nature of shorter-term trading efficacy.
To demonstrate the advantages of the Transformer factor over traditional factors, 100 price-volume type factors from JoinQuant  [^3] https://www.joinquant.com/ (one of Chinese largest financial quantitative platforms) were selected for backtesting under the same trading strategy (Algorithm 1). Price-volume factors were used because these are calculated based on stocks’ prices, volumes, and turnover rates, which are similar to the training data used by Transformer factors (also used stock turnover rates and price data). The detailed performance of factors and benchmark is shown in Appendix A. On average, these factors achieved an annual return of -3.78%, an average excess return of -2.15%, and a Sharpe ratio of -0.36, where the benchmark return is -1.77%. The average maximum drawdown of the strategy was 44.88%, which means the potential worst-case scenario or the most extreme possible loss. The average volatility of the strategy ($\sigma_{p}$) was 0.26, with the benchmark at 0.197.
Figure 4: Backtest results based on different factors
Among these factor-based strategies, the Transformer factor performed well. Except for the win rate, which was slightly lower than a few strategies, it excelled in terms of annual return, annual excess return, $\sigma_{p}$, and Sharpe Ratio, ranking the best among the 101 factors. Figure 4 illustrates the return curves of the Transformer (Trans_Month_1) alongside the benchmark and some traditional factors. ”25_week_rank” represents the current price’s position over the past 25 weeks; ”EMAC20” and ”MAC20” are the 20-day index moving average and the stock’s 20-day moving average respectively; ”DAVOL20” and ”VOL20” represent the ratio of the 20-day average turnover rate to the 120-day average turnover rate, and the mean of the stock’s 20-day turnover rate; ”ROC20” is the price rate of change over 20 days; ”MAWVAD” is calculated as the product of the difference between the closing price and the opening price, divided by the range of the highest and lowest prices, all multiplied by the volume, accumulated over six days; ”Variance20” is the variance of the stock’s 20-day annualized returns. Most of these factors are compared based on a window of 20 timestamps, as the Transformer factor was also trained on 20 timestamps. In comparison with these factors, the Transformer factor (blue line) demonstrated a significantly better performance in terms of returns than the benchmark (orange line) and other factor strategies, expressing the improvement of the Transformer factor over traditional price-volume factors.
5.2 Half-life period test
To test the half-life period of the factor, another three portfolios based on $Month\_3$ were conducted. The first one is the stock pool used in the next month; the second is using it in the month after the month; the last one is using it in the third month after the last feature of inputs. For instance, in the first strategy, the trade will start in January 2020; in the second, it will start in February 2020; in the second, it will start in March 2020. For the period from January 2020 to February 2022, the strategy achieved a total return of 65.29% and an annual return of 17.87%. The excess return over this period was 66.39%, indicating a strong performance above a benchmark or risk-free rate. The Sharpe Ratio for this period was 0.863, which is a respectable figure that suggests a favorable risk-adjusted return. Additionally, the strategy generated an alpha of 0.164, demonstrating its ability to outperform the expected return on a risk-adjusted basis.
Table 3: Result of the same monthly strategy with different start month
Strategy Total return Annual return Excess return Sharpe ratio Alpha
2020/01-2022/02 65.29% 17.87% 66.39% 0.863 0.164
2020/02-2022/03 67.52% 18.21% 65.58% 0.940 0.161
2020/03-2022/04 56.41% 15.63% 52.96% 0.739 0.135
In the subsequent period from February 2020 to March 2023, the strategy yielded a slightly higher total return of 67.52% and an annual return of 18.21%. The excess return was 65.58%, with an improved Sharpe Ratio of 0.940, reflecting a better risk-adjusted return compared to the previous period. The alpha for this interval was 0.161, indicating the strategy’s consistent performance in delivering additional returns.
Lastly, Table 3 shows the result of the backtest. From March 2020 to April 2023, the strategy’s total return decreased to 56.41%, with an annual return of 15.63% and an excess return of 52.96%. The Sharpe Ratio also declined to 0.739, but it still represents a relatively efficient investment compared to the volatility of returns. The alpha reduced to 0.135, which, while lower than the other periods, still indicates that the strategy managed to provide returns above those predicted by its risk profile.
5.3 Discussion
Comparing the results between monthly, weekly, and daily strategies provides further insights. The monthly strategies can deliver the highest returns, suggesting that longer-term signals better capture short-term trends. The weekly strategies underperform monthly ones but outperform the daily frequency overall. This indicates weekly data strikes a balance between filtering noise and responding swiftly to emerging patterns.
The daily frequency strategies perform lower annual returns compared to others. This highlights the challenge of making profitable trades at higher frequencies - noise and volatility make reliable signals more difficult. The distribution of returns also shows increased dispersion at the daily level. However, daily strategies achieve more consistent turnover rates around 5-7%. This allows for portfolio adjustments while avoiding excessive trading costs. In contrast, monthly strategies see a turnover rate above 50% in some cases. There is likely an optimal point between 1 week to 1 month where matches market regularly.
6 Conclusion
This study proposed an approach to exploit transformer models for quantitative stock prediction and trading. We addressed the need for handling numerical input data rather than text and adapting the model for forecasting tasks rather than sequence-to-sequence problems common in NLP. To enable direct processing of numerical time series data, we replaced the word embedding layer with a standard linear layer and removed the output masking operations. We also simplified the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
Eventually, we tested the factor in the reality back test of the financial market and compared the factor with other 100 price-volume factors. Our experimental results demonstrate the promise of this approach. The model-based trading strategies were able to deliver substantial excess returns over the benchmark under a longer time period, with Sharpe Ratios indicating sound risk-adjusted performance. The positive alphas affirm the strategy’s ability to outperform expected returns after accounting for risk factors.
There is further scope to enhance the framework, for instance by incorporating additional signals like news and fundamentals as auxiliary inputs. The self-attention mechanism could likely be improved to encode greater temporal context. Or use transformer-based models such as GPT-4 [45] and Claude3 [46] to establish the strategy as these models are fine-tuned and may perform better in quantitative finance tasks.
Overall, our work illustrates the viability of transformers for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables. With further research to solve the above limitations, we believe such ML-based quantitative methods hold rich potential. The framework provides a flexible foundation for better understanding markets and developing profitable trading strategies.
7 Acknowledgements and declaration
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix A Detailed backtest results between 100 factors and the transformer factor
All the data of factors comes from JoinQuant [^4] https://www.joinquant.com/, which is one of the most common quantitative finance platforms. Detailed descriptions of the following factors are at JoinQuant Factors.
Seven metrics are used to evaluate the performance of each factor:
  • [label=•]
  • Annual Return (AR) and Annual Excess Return (AER) are two metrics to reflect the return of the portfolio directly. The excess returns are returns achieved above and beyond the return of a proxy and the CSI300 index is used as the basis when calculating the excess returns.
  • Win Rate (WR) is the metric to evaluate the fraction between the action that makes a profit and all the actions.
  • Sharpe Ratio (SR) is a measure of risk-adjusted return that can describe the additional earnings an investor receives per unit of increase, which is shown in Eq.(8).
  • The Alpha measures the portfolio in generating returns above the market benchmark. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment.
  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to the change in the market.
  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss.
  • Portfolio volatility ($\sigma_{p}$) is used to measure the risk of a strategy, where is positively correlated to the risk of a strategy.
Factor AR (%) AER (%) WR (%) SR Alpha Beta MD (%) $\sigma_{p}$
ARBR -1.47 0.32 54.7 -0.251 -0.002 0.914 34.46 0.221
AR -5.87 -4.37 53.7 -0.406 -0.048 0.981 47.84 0.260
ATR14 -2.12 -0.37 62.7 -0.250 0.006 1.160 46.55 0.250
ATR6 -3.94 -2.31 61.2 -0.328 -0.014 1.175 50.11 0.253
BBIC 1.17 3.14 50.9 -0.109 0.033 1.038 32.64 0.258
BR -6.91 -5.47 53.6 -0.467 -0.064 0.934 51.72 0.254
CCI10 -18.65 -17.99 43.3 -1.402 -0.261 0.812 72.04 0.220
CCI15 -17.70 -16.98 44.2 -1.293 -0.240 0.832 71.43 0.223
CCI20 -15.52 -14.66 46.3 -1.051 -0.196 0.868 69.12 0.235
CR20 -4.27 -2.67 54.8 -0.329 -0.032 0.927 49.92 0.263
DAVOL10 -7.44 -6.04 54.1 -0.504 -0.071 0.920 45.45 0.248
DAVOL20 -10.95 -9.78 54.4 -0.702 -0.119 0.919 48.96 0.246
DAVOL5 -9.54 -8.28 51.8 -0.612 -0.097 0.950 46.41 0.250
EMA5 2.77 4.85 48.3 -0.047 0.049 1.041 30.70 0.259
EMAC10 0.27 2.17 50.6 -0.144 0.024 1.037 31.42 0.258
EMAC12 1.58 3.57 50.6 -0.093 0.037 1.032 31.76 0.258
EMAC20 6.01 8.29 54.1 0.069 0.079 1.044 32.43 0.260
EMAC26 3.83 5.97 54.1 -0.008 0.059 1.046 34.87 0.261
EMAC120 1.74 3.75 48.3 -0.088 0.034 0.960 35.81 0.254
Kurtosis20 -2.96 -1.27 52.5 -0.395 -0.027 0.772 27.71 0.182
Kurtosis60 -0.28 1.59 53.9 -0.240 -0.001 0.716 22.07 0.179
Kurtosis120 -4.12 -2.51 49.0 -0.468 -0.043 0.708 28.28 0.182
MAC5 0.54 2.46 46.9 -0.134 0.026 1.035 30.03 0.257
MAC10 -1.20 0.61 49.6 -0.206 0.008 1.039 34.15 0.256
MAC20 2.47 4.52 53.5 -0.059 0.046 1.043 34.58 0.257
MAC60 11.10 13.72 53.8 0.243 0.120 0.987 28.61 0.254
MAC120 3.11 5.21 47.7 -0.035 0.049 0.987 35.40 0.259
MACDC -11.12 -9.97 52.4 -0.674 -0.117 1.003 46.48 0.261
MASS -10.55 -9.35 51.0 -0.673 -0.115 0.886 48.34 0.249
MAWVAD 7.29 9.67 55.9 0.119 0.082 0.891 36.54 0.248
MFI14 -6.21 -4.73 51.5 -0.456 -0.056 0.918 45.65 0.241
PLRC6 -14.83 -13.92 47.8 -0.905 -0.178 0.960 66.24 0.259
PLRC12 -7.28 -5.87 53.3 -0.452 -0.064 1.012 52.30 0.273
PLRC24 -5.77 -4.26 58.1 -0.361 -0.044 1.034 53.28 0.289
PSY -9.11 -7.83 50.6 -0.660 -0.096 0.880 49.79 0.223
Price1M -13.55 -12.55 53.0 -0.760 -0.156 0.977 64.85 0.280
Price3M -6.69 -5.25 58.4 -0.380 -0.056 1.023 59.14 0.305
Price1Y 1.55 3.54 66.8 -0.075 0.041 1.116 60.25 0.326
ROC6 -16.13 -15.30 47.4 -0.982 -0.200 0.978 68.14 0.262
ROC20 -11.63 -10.51 54.7 -0.628 -0.123 1.014 60.28 0.291
ROC60 -6.86 -5.43 59.6 -0.382 -0.056 1.062 62.16 0.309
ROC120 2.36 4.41 66.9 -0.051 0.049 1.117 56.36 0.319
Skewness20 -5.91 -4.41 50.8 -0.538 -0.058 0.820 40.65 0.197
Skewness60 -6.47 -5.01 51.9 -0.595 -0.068 0.765 38.02 0.190
Skewness120 -4.37 -2.77 51.1 -0.459 0.044 0.743 36.01 0.192
TRIX5 -10.72 -9.54 52.5 -0.607 -0.110 1.009 56.07 0.280
TRIX10 -3.51 -1.85 58.4 -0.268 -0.019 1.008 52.92 0.292
TVMA20 -1.92 -0.16 64.9 -0.209 0.014 1.268 44.66 0.288
TVMA6 -0.75 1.09 62.4 -0.166 0.026 1.259 39.46 0.288
TVSTD20 -1.83 -0.06 58.5 -0.214 0.012 1.218 35.89 0.277
TVSTD6 -3.77 -2.13 53.4 -0.297 -0.010 1.202 46.57 0.272
VDEA -3.05 -1.37 57.6 -0.305 -0.023 0.857 37.01 0.239
VDIFF -7.93 -6.57 0.2 -0.533 -0.079 0.895 39.79 0.247
VEMA5 -0.41 1.46 57.9 -0.186 0.006 0.850 37.31 0.237
VEMA10 0.95 2.91 58.9 -0.128 0.019 0.844 35.51 0.236
VEMA12 1.01 2.96 59.3 -0.126 0.020 0.840 35.17 0.235
VEMA26 2.41 4.46 59.3 -0.068 0.033 0.838 31.03 0.231
VMACD -1.03 0.79 50.5 -0.216 0.001 0.875 33.96 0.235
VOL5 -8.12 -6.76 55.9 -0.452 -0.066 1.155 53.86 0.296
VOL10 -7.07 -5.64 58.3 -0.404 -0.053 1.145 50.65 0.298
VOL20 -5.43 -3.89 58.6 -0.337 -0.032 1.156 50.19 0.298
VOL60 0.42 2.34 58.1 -0.125 0.032 1.145 38.35 0.286
VOL120 -1.88 -0.12 53.0 -0.213 0.007 1.136 42.09 0.281
VOL240 -5.51 -3.98 47.0 -0.382 -0.038 1.075 43.53 0.265
VOSC -6.35 -4.88 51.2 -0.453 -0.056 0.944 40.40 0.246
VROC6 -16.30 -15.49 42.7 -1.179 -0.208 0.897 69.03 0.221
VROC12 -11.47 -10.34 46.7 -0.793 -0.127 0.915 51.62 0.228
VR -8.43 -7.10 52.8 -0.566 -0.083 0.932 48.57 0.244
VSTD10 -0.77 1.07 56.3 -0.205 0.001 0.843 37.75 0.234
VSTD20 -1.60 0.19 53.7 -0.245 -0.007 0.844 37.31 0.232
Variance20 -1.94 -0.18 59.9 -0.201 0.005 1.124 46.38 0.302
Variance60 9.96 12.51 65.2 0.168 0.122 1.178 30.63 0.313
Variance120 5.40 7.65 60.8 0.040 0.083 1.193 39.54 0.312
Volume1M -15.54 -14.68 20.7 -0.867 -0.189 0.991 63.68 0.285
WVAD -0.71 1.13 52.3 -0.199 0.002 0.836 42.38 0.238
arron_down_25
-7.35 -5.94 48.3 -0.593 -0.072 0.880 42.63 0.209
arron_up_25 -15.51 -14.65 46.4 -1.049 -0.195 0.870 66.73 0.235
bear_power -13.31 -12.30 50.8 -0.800 -0.154 0.934 59.58 0.261
beta 10.30 12.87 62.4 0.168 0.134 1.344 35.96 0.328
boll-up 2.19 4.23 54.5 -0.068 0.044 1.062 33.09 0.265
book_to_price 1.87 3.86 58.5 -0.117 0.031 0.889 26.73 0.181
bull_power -14.81 -13.91 51.1 -0.844 -0.176 0.984 65.99 0.278
earnings_yield -2.41 -0.68 52.6 -0.348 -0.029 0.627 25.87 0.188
growth 2.90 4.99 59.9 -0.045 0.052 1.068 37.35 0.244
leverage 2.05 4.07 53.1 -0.089 0.023 0.719 22.57 0.217
liquidity -2.71 -1.00 56.7 -0.243 -0.002 1.137 41.28 0.284
momentum 10.93 13.54 65.8 0.195 0.129 1.159 43.19 0.310
money_flow_20 -1.72 0.05 65.5 -0.202 0.016 1.269 44.10 0.288
price_no_fq 6.83 9.17 76.1 0.095 0.092 1.132 44.86 0.267
pull-down -2.28 -0.54 47.6 -0.317 -0.015 0.831 34.46 0.203
residual_volatility
4.13 6.30 62.3 0.003 0.062 1.048 41.34 0.269
sharpe_ratio_20
-11.88 -10.78 54.0 -0.663 -0.129 0.979 57.66 0.282
sharpe_ratio_60
-7.42 -6.02 58.8 -0.417 -0.065 1.029 65.19 0.300
sharpe_ratio_120
-0.32 1.55 65.7 -0.143 0.019 1.062 60.52 0.303
size -6.36 -4.89 51.8 -0.571 -0.058 0.906 36.00 0.196
turnover_volatility
-10.80 -9.62 51.8 -0.624 -0.108 1.075 57.05 0.274
1day_VPT -16.45 -15.65 44.8 -1.062 -0.211 0.904 70.43 0.248
1day_VPT_6
-6.36 -4.89 53.0 -0.416 -0.053 0.995 47.06 0.269
1day_VPT_12
-0.07 1.82 57.9 -0.146 0.020 1.033 46.51 0.278
25week_close
-11.51 -10.39 51.0 -0.859 -0.133 0.817 47.79 0.211
Trans_Month_1
17.35 19.43 57.8 0.915 0.162 0.588 18.35 0.161
Average -3.78 -2.15 53.71 -0.36 -0.03 0.98 44.88 0.26
CSI300 1.77 \ \ \ \ \ \ 0.197